# Parametric Scaling Law of Tuning Bias in Conformal Prediction

### v1

@@ Appendix C Some Useful Lemmas and Corollaries
The ConfTr(op) is a regularization case of ConfTr, where the class $\mathcal{H}_{\text{op}}$ is used as the constraint. Since the class $\mathcal{H}_{\text{op}}$ defined in Eq. (14) is a subset of $\mathcal{H}$, we have
$\mathbb{E}\sup_{f\in\mathcal{H}_{\text{op}}}\Big{|}\frac{1}{n}\sum_{i\in[n]}f(\boldsymbol{x}_{i},y_{i})-\mathbb{E}[f(\boldsymbol{x}_{\text{test}},y_{\text{test}})\mid\mathcal{D}_{\text{cal}}]\Big{|}\leq\mathbb{E}\sup_{f\in\mathcal{H}}\Big{|}\frac{1}{n}\sum_{i\in[n]}f(\boldsymbol{x}_{i},y_{i})-\mathbb{E}[f(\boldsymbol{x}_{\text{test}},y_{\text{test}})\mid\mathcal{D}_{\text{cal}}]\Big{|},$
@@
$\mathop{TuningBias}(\widehat{\mathcal{C}}_{\text{ConfTr(op)}})\leq\mathop{TuningBias}(\widehat{\mathcal{C}}_{\text{ConfTr}}).$
Appendix K Additional Theoretical Results of Score aggregation
For a more general case about the selection of scores, we could also apply the same analysis to the score aggregation (Luo & Zhou, 2024). In the setting of score aggregation (Luo & Zhou, 2024), they choose an aggregation weights vector $\boldsymbol{w}$ to aggregate multiple score functions:

### v2

@@ Appendix D Some Useful Lemmas and Corollaries
The ConfTr(op) is a regularization case of ConfTr, where the class $\mathcal{H}_{\Lambda,\text{op}}$ is used as the constraint. Since the class $\mathcal{H}_{\Lambda,\text{op}}$ defined in Eq. (14) is a subset of $\mathcal{H}_{\Lambda}$, we have
$\mathbb{E}\sup_{f\in\mathcal{H}_{\Lambda,\text{op}}}\Big{|}\frac{1}{n}\sum_{i\in[n]}f(\boldsymbol{x}_{i},y_{i})-\mathbb{E}[f(\boldsymbol{x}_{\text{test}},y_{\text{test}})\mid\mathcal{D}_{\text{cal}}]\Big{|}\leq\mathbb{E}\sup_{f\in\mathcal{H}_{\Lambda}}\Big{|}\frac{1}{n}\sum_{i\in[n]}f(\boldsymbol{x}_{i},y_{i})-\mathbb{E}[f(\boldsymbol{x}_{\text{test}},y_{\text{test}})\mid\mathcal{D}_{\text{cal}}]\Big{|},$
@@
$\mathop{\text{TuningBias}}({\mathcal{C}}_{\text{ConfTr (op)}})\leq\mathop{\text{TuningBias}}({\mathcal{C}}_{\text{ConfTr}}).$
Order-preserving regularization for vector scaling
For the vector scaling defined in Eq. (8), because the softmax function is order-preserving, the order-preserving regularization on the output probability in Eq. (7) is equivalent to the order-preserving regularization on the input logits $f(\boldsymbol{x})$. If we apply the order-preserving regularization on the vector scaling, we should have for any logits vector $f(\boldsymbol{x})$,
$\forall j,j^{\prime}\in[K],f_{j}(\boldsymbol{x})\leq f_{j^{\prime}}(\boldsymbol{x})\iff W_{j}f_{j}(\boldsymbol{x})+b_{j}\leq W_{j^{\prime}}f_{j^{\prime}}(\boldsymbol{x})+b_{j^{\prime}},$
where $f_{j}(\boldsymbol{x})$ is the $j$-th element of the logits vector $f(\boldsymbol{x})$, $W_{j}$ is the $j$-th element of $W$, and $b_{j}$ is the $j$-th element of $b$.
Lemma K.1.
Vector scaling is order-preserving if and only if $W_{j}=W_{j^{\prime}}>0$ for all $j,j^{\prime}\in[K]$ and $b_{j}=b_{j^{\prime}}$ for all $j,j^{\prime}\in[K]$.
Proof.
()($\Leftarrow$)( ⇐ ) It is a direct result of the definition of vector scaling.
()($\Rightarrow$)( ⇒ ) If vector scaling is order-preserving, then
$\forall j,j^{\prime}\in[K],f_{j}(\boldsymbol{x})\leq f_{j^{\prime}}(\boldsymbol{x})\iff W_{j}f_{j}(\boldsymbol{x})+b_{j}\leq W_{j^{\prime}}f_{j^{\prime}}(\boldsymbol{x})+b_{j^{\prime}}.$
We first exclude the case when $W_{j}<0$ for any $j\in[K]$: For any $W_{j}<0$, and fixed $b_{j},b_{j^{\prime}}$, we could find small enough $f_{j}(\boldsymbol{x})$, such that $f_{j}(\boldsymbol{x})<f_{j^{\prime}}(\boldsymbol{x})$ and $W_{j}f_{j}(\boldsymbol{x})+b_{j}>W_{j^{\prime}}f_{j^{\prime}}(\boldsymbol{x})+b_{j^{\prime}}$, which contradicts the order-preserving property. Therefore, we only consider the case when $W_{j},W_{j^{\prime}}>0$ for all $j,j^{\prime}\in[K]$. Without loss of generality, we assume $f_{j}(\boldsymbol{x})\leq f_{j^{\prime}}(\boldsymbol{x})$ and $W_{j}<W_{j^{\prime}}$, and let $f_{j^{\prime}}(\boldsymbol{x})=f_{j}(\boldsymbol{x})+\epsilon$ with $\epsilon\geq 0$, then we have
$W_{j}f_{j}(\boldsymbol{x})+b_{j}\leq W_{j^{\prime}}f_{j^{\prime}}(\boldsymbol{x})+b_{j^{\prime}}\iff(W_{j}-W_{j^{\prime}})f_{j}(\boldsymbol{x})\leq b_{j^{\prime}}-b_{j}+W_{j^{\prime}}\epsilon$
Let $\varepsilon\to 0$, we have
$(W_{j}-W_{j^{\prime}})f_{j}(\boldsymbol{x})\leq b_{j^{\prime}}-b_{j}$
hold for any $f_{j}(\boldsymbol{x})$. We could find a small enough $f_{j}(\boldsymbol{x})$, such that
$(W_{j}-W_{j^{\prime}})f_{j}(\boldsymbol{x})>b_{j^{\prime}}-b_{j}$
This contradicts the order-preserving property. Therefore, we must have $W_{j}=W_{j^{\prime}}>0$ for all $j,j^{\prime}\in[K]$ and $b_{j}=b_{j^{\prime}}$ for all $j,j^{\prime}\in[K]$. ∎
Remark K.2.
By Proposition 5.2, if we apply the order-preserving regularization on the vector scaling, the parameter space of the vector scaling is reduced to $\Lambda=\mathbb{R}^{2}$.
Order-preserving regularization for matrix scaling
Here we consider a more complex case that the parameter space of the scaling is a matrix: $Wf(\boldsymbol{x})+b$, where $f$ be a logits value function for a classifier of $K$ classes, and $W$ is a matrix of size $K\times K$ and $b$ is a vector of size $K$. Here the dimension of the parameter space is $K^{2}+K$. We apply the order-preserving regularization on the matrix scaling:
Lemma K.3.
Let $f$ be a logits value function for a classification of $K$ classes. The matrix scaling $g(x)=Wf(x)+b$ is order-preserving if and only if $W$ has the form $W=aI+\mathbf{1}v^{T}$ for some scalar $a>0$ and vector $v\in\mathbb{R}^{K}$, and $b$ is a constant vector (i.e., $b_{j}=b_{j^{\prime}}$ for all $j,j^{\prime}\in[K]$). Here, $I$ is the $K\times K$ identity matrix and $\mathbf{1}$ is the $K$-dimensional vector of all ones.
Remark K.4.
Here, we regard C-Adapter as a special case of matrix scaling with order-preserving regularization. The above proposition shows that the order-preserving regularization reduces the dimension of parameter space from $K^{2}+K$ to $K+2$, which is much smaller than VS with its dimension being $2K$. Based on the parametric scaling law (Section 4), we explain why C-Adapter can achieve lower tuning bias than VS. And Lemma K.1 is a special case of Lemma K.3.
Appendix L Additional Theoretical Results of Score aggregation
For a more general case about the selection of scores, we could also apply the same analysis to the score aggregation (Luo & Zhou, 2024). In the setting of score aggregation, they choose an aggregation weights vector $\boldsymbol{w}$ to aggregate multiple score functions: